# Existence and non-existence of consistent estimators in supercritical controlled branching processes

### v1

@@ 5 Proofs
Let $X,X_{1},\dots,X_{n}$, $n\in\mathbbm{N}_{1}$, be i.i.d. random variables on $\mathbbm{N}_{0}$ with $\mathbbm{E}X:=m$, $\text{Var}(X):=\sigma^{2}>0$, and finite third absolute central moment $\rho:=\mathbbm{E}|X-m|^{3}$. Define $S_{n}:=\sum_{i=1}^{n}X_{i}$, and let $W_{n}$ be a random variable with a $\text{DN}(nm,m\sigma^{2})$ distribution. Then
$||\mathcal{L}_{S_{n}}-\mathcal{L}_{W_{n}}||_{TV}$ $\leq\sqrt{\frac{2}{\pi}}\left(\frac{3\rho}{\sigma^{2}}+2\right)\Big{(}1+4(n-1)\left(1-||\mathcal{L}_{X}-\mathcal{L}_{(X+1)}||_{TV}\right)\Big{)}^{-\frac{1}{2}}$
Proof. We can use [4, Theorem 7.4], simplified to the case of an i.i.d. sum, to see that
$||\mathcal{L}_{S_{n}}-\mathcal{L}_{W_{n}}||_{TV}\leq\left(\frac{3\rho}{2\sigma^{2}}+1\right)||\mathcal{L}_{S_{n-1}}-\mathcal{L}_{(S_{n-1}+1)}||_{TV}+\left(5+3\sqrt{\frac{\pi}{8}}\right)\frac{\rho}{\sqrt{n}\sigma^{3}}+\frac{1}{2\sqrt{2\pi n}\sigma}.$
A bound on $||\mathcal{L}_{S_{n-1}}-\mathcal{L}_{(S_{n-1}+1)}||_{TV}$ is then provided by [25, Corollary 1.6], as
$||\mathcal{L}_{S_{n-1}}-\mathcal{L}_{(S_{n-1}+1)}||_{TV}$ $\leq\sqrt{\frac{2}{\pi}}\left(\frac{1}{4}+(n-1)\left(1-||\mathcal{L}_{X}-\mathcal{L}_{(X+1)}||_{TV}\right)\right)^{-\frac{1}{2}}$
$=2\sqrt{\frac{2}{\pi}}\Big{(}1+4(n-1)\left(1-||\mathcal{L}_{X}-\mathcal{L}_{(X+1)}||_{TV}\right)\Big{)}^{-\frac{1}{2}}.$
Put together, this yields the desired result. $\blacksquare$
Lemma 5.6.
$\mathbbm{E}|X-m|^{3}\leq 8(\gamma+3m\sigma^{2}+m^{3}).$
Proof. By Minkowski’s inequality, $(\mathbbm{E}|X-m|^{3})^{1/3}\leq(E|X|^{3})^{1/3}+(E|-m|^{3})^{1/3}=(E|X|^{3})^{1/3}+|m|$. Using Jensen’s inequality and subsequently Hölder’s inequality, we see that $|m|\leq\mathbbm{E}|X|\leq(E|X|^{3})^{1/3}$. It follows that $(\mathbbm{E}|X-m|^{3})^{1/3}\leq 2(E|X|^{3})^{1/3}$, from which we can conclude $\mathbbm{E}|X-m|^{3}\leq 8\cdot\mathbbm{E}|X|^{3}$.
However, since $X$ is supported on $\mathbbm{N}_{0}$, $\mathbbm{E}|X|^{3}=\mathbbm{E}X^{3}$, and we can show by expanding $\mathbbm{E}(X-m)^{3}$ that $\mathbbm{E}X^{3}=\gamma+3m\sigma^{2}+m^{3}$. The desired result follows. $\blacksquare$
Lemma 5.7.
$||\mathcal{L}_{W}-\mathcal{L}_{\tilde{W}}||_{TV}\leq\frac{3|\sigma^{2}-\tilde{\sigma}^{2}|}{2(\sigma^{2}\vee\tilde{\sigma}^{2})}+\frac{|m-\tilde{m}|}{2(\sigma\vee\tilde{\sigma})}.$
Proof. We can find
$||\mathcal{L}_{W}-\mathcal{L}_{\tilde{W}}||_{TV}$ $=\frac{1}{2}\sum_{n\in\mathbbm{Z}}\big{|}\mathbbm{P}(W=n)-\mathbbm{P}(\tilde{W}=n)\big{|}$
[…]
$=||N(m,\sigma^{2})-N(\tilde{m},\tilde{\sigma}^{2})||_{TV}.$
$||N(m,\sigma^{2})-N(\tilde{m},\tilde{\sigma}^{2})||_{TV}\leq\frac{3|\sigma^{2}-\tilde{\sigma}^{2}|}{2(\sigma^{2}\vee\tilde{\sigma}^{2})}+\frac{|m-\tilde{m}|}{2(\sigma\vee\tilde{\sigma})}.$
$\blacksquare$
Having completed this initial set-up, we are ready to prove Lemma 3.7.
Proof of Lemma 3.7. Suppose that $\{Z_{n},z_{0}\},\{X_{n},z_{0}\}\in\Pi^{(u)}$, such that their respective control functions $\phi_{Z}(\cdot)$ and $\phi_{X}(\cdot)$ are both linearly-divisible and satisfy (8), their offspring distributions $\xi_{Z}$ and $\xi_{X}$ both have finite third moments and lattice size one, and that there exists $r<1$ such that $|\mathbbm{E}(Z_{1}|Z_{0}=z)-\mathbbm{E}(X_{1}|X_{0}=z)|=O(z^{r/2})$ and $|\text{Var}(Z_{1}|Z_{0}=z)-\text{Var}(X_{1}|X_{0}=z)|=O(z^{r})$.
$\tilde{\xi}_{Z}(z):=\sum_{i=1}^{\tilde{\phi}_{Z}(z)}\xi_{Z,i},$
while
$\tilde{\gamma}_{Z}(z)=\frac{\varepsilon_{Z}(z)\cdot\gamma_{Z}+\upsilon_{Z}(z)\cdot m_{Z}^{3}+\nu_{Z}^{2}(z)\cdot m_{Z}\cdot\sigma_{Z}^{2}}{z}\cdot\frac{z}{l_{Z}(z)}.$
$\tilde{m}_{Z}(z)\leq\lambda_{Z}a_{Z}m_{Z},\quad\tilde{\sigma}_{Z}^{2}(z)\leq\lambda_{Z}(a_{Z}\sigma_{Z}^{2}+b_{Z}m_{Z}^{2}),\quad\text{and}\quad\tilde{\gamma}_{Z}(z)\leq\lambda_{Z}(a_{Z}\gamma_{Z}+c_{Z}m_{Z}^{3}+b_{Z}m_{Z}\sigma_{Z}^{2}).$ (25)
In addition:

### v2

@@ 6 Proofs of consistency results
Let $\{Z_{n},z_{0}\}$ be a supercritical CBP with control function $\phi(\cdot)$ having $\varepsilon(z)=\alpha z$ and $\nu^{2}(z)=\beta z$, and with $\sigma^{2}$ finite. Then, on the event $\{Z_{n}\to\infty\}$ (so that $\phi(Z_{n})\to\infty$ as well) and for $s$ such that $1<s<m\alpha$,
$\lim_{n\to\infty}\mathbbm{P}\bigg{(}\sum_{k=1}^{n}\frac{\phi(Z_{k-1})}{\phi(Z_{n-1})}\leq\frac{1}{1-s^{-1}}\bigg{)}=1.$
Proof of Theorem 3.13:.
Let $\{Z_{n},z_{0}\}$ be a supercritical CBP with unknown control function $\phi(\cdot)$ having $\varepsilon(z)=\alpha z$ and $\nu^{2}(z)=\beta z$, and with $\sigma^{2}$ finite. Assume that progenitor counts are observed alongside population numbers, and accordingly take $\mathcal{F}_{n}$ to be the $\sigma$-algebra generated by $\big{(}Z_{0},\,\phi(Z_{0}),\dots,\phi(Z_{n-1}),\,Z_{n}\big{)}$ and $\mathcal{F}_{\phi_{n}}$ the $\sigma$-algebra generated by $\big{(}Z_{0},\,\phi(Z_{0}),\dots,Z_{n},\,\phi(Z_{n})\big{)}$, while $\mathcal{F}_{\phi_{-1}}$ is taken to represent the trivial $\sigma$-algebra.
Throughout the proof, we make the simplifying assumption that $\mathbbm{P}(\xi>0)=1$ and $\mathbbm{P}(\phi(z)>0)=1$ for all $z\in\mathbbm{N}_{1}$. This implies that $\mathbbm{P}(Z_{1}>0\,|\,Z_{0}=z)=1$ for all $z\in\mathbbm{N}_{1}$, that $I^{\phi}_{n}=\{1,\dots,n\}$, and that $I^{-}_{n}=\{0,\dots,n-1\}$. Additionally, under this assumption, $\mathbbm{P}(Z_{n}\to\infty)=1$, and the estimators simplify to
$\hat{m}_{n}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{Z_{k}}{\phi(Z_{k-1})},$ $\quad\hat{\alpha}_{n}$ $:=\frac{1}{n}\sum_{k=0}^{n-1}\frac{\phi(Z_{k})}{Z_{k}},$
$\bar{\sigma}_{n}^{2}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-m\cdot\phi(Z_{k-1}))^{2}}{\phi(Z_{k-1})},$ $\quad\bar{\beta}_{n}$ $:=\frac{1}{n}\sum_{k=0}^{n-1}\frac{(\phi(Z_{k})-\alpha\cdot Z_{k})^{2}}{Z_{k}},$
$\hat{\sigma}_{n}^{2}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-\tilde{m}_{n}\cdot\phi(Z_{k-1}))^{2}}{\phi(Z_{k-1})},$ $\quad\hat{\beta}_{n}$ $:=\frac{1}{n}\sum_{k=0}^{n-1}\frac{(\phi(Z_{k})-\tilde{\alpha}_{n-1}\cdot Z_{k})^{2}}{Z_{k}}.$
We omit the extension of our proof to the case where $\mathbbm{P}(\xi=0)>0$ and/or there exists $z\in\mathbbm{N}_{1}$ such that $\mathbbm{P}(\phi(z)=0)>0$, which follows similar arguments as in the proof of Theorem 3.5, part (i), case (b).
Throughout the rest of the proof, we will also make use of the following:
$\mathbbm{E}\big{(}\phi^{-1}(z)\big{)}\leq 1\quad\text{for all}\;z\in\mathbbm{N}_{1}.$ (34)
This follows as a consequence of the assumption $\mathbbm{P}(\phi(z)>0)=1$ for all $z\in\mathbbm{N}_{1}$, which tells us that $\phi(z)\geq 1$ a.s., and therefore $\phi^{-1}(z)\leq 1$ a.s..
(i) Strong consistency of $\boldsymbol{\hat{m}_{n}}$: Let $U_{n}:=\tilde{m}_{n}-m$, where we recall $\tilde{m}_{n}:=Z_{n}/\phi(Z_{n-1})$, so that
$\mathbbm{E}(U_{n}|\mathcal{F}_{n-1})=\mathbbm{E}\big{(}\mathbbm{E}(U_{n}|\mathcal{F}_{\phi_{n-1}})\big{|}\mathcal{F}_{n-1}\big{)}=0$
and $\mathbbm{E}|U_{n}|\leq 2m$, while, using (34),
$\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})=\mathbbm{E}\big{(}\text{Var}(U_{n}|\mathcal{F}_{\phi_{n-1}})\big{|}\mathcal{F}_{n-1}\big{)}=\mathbbm{E}\bigg{(}\frac{\sigma^{2}}{\phi(Z_{n-1})}\bigg{|}\mathcal{F}_{n-1}\bigg{)}\leq\sigma^{2}.$
Therefore
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})<\infty,$
so that, by Corollary 6.2, $\frac{1}{n}\sum_{k=1}^{n}U_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\hat{m}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}m$.
(ii) Strong consistency of $\boldsymbol{\hat{\alpha}_{n}}$: Let $U_{\phi_{n}}:=\tilde{\alpha}_{n}-\alpha$, where we recall $\tilde{\alpha}_{n}:=\phi(Z_{n})/Z_{n}$, so that
$\mathbbm{E}(U_{\phi_{n}}|\mathcal{F}_{\phi_{n-1}})=\mathbbm{E}\big{(}\mathbbm{E}(U_{\phi_{n}}|\mathcal{F}_{n})\big{|}\mathcal{F}_{\phi_{n-1}}\big{)}=0$
and $\mathbbm{E}|U_{\phi_{n}}|\leq 2\alpha$, while
$\mathbbm{E}\big{(}U_{\phi_{n}}^{2}\big{|}\mathcal{F}_{\phi_{n-1}}\big{)}=\mathbbm{E}\big{(}\text{Var}(U_{\phi_{n}}|\mathcal{F}_{n})\big{|}\mathcal{F}_{\phi_{n-1}}\big{)}=\mathbbm{E}\bigg{(}\frac{\beta}{Z_{n}}\bigg{|}\mathcal{F}_{n}\bigg{)}\leq\beta$
since $Z_{n}\geq 1$ under our simplifying assumption that $\mathbbm{P}(\phi(z)=0)=0$ and $\mathbbm{P}(\xi=0)=0$. Therefore
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}\big{(}U_{\phi_{n-1}}^{2}\big{|}\mathcal{F}_{\phi_{n-2}}\big{)}<\infty,$
so that, by Corollary 6.2, $\frac{1}{n}\sum_{k=1}^{n}U_{k-1}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\hat{\alpha}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\alpha$.
(iii) Strong consistency of $\boldsymbol{\bar{\sigma}^{2}_{n}}$: Under the assumption that $m$ is known, we introduce the estimator
$\tilde{\sigma}^{2}_{n}:=\frac{(Z_{n}-m\cdot\phi(Z_{n-1}))^{2}}{\phi(Z_{n-1})},$
such that $\bar{\sigma}^{2}_{n}=\frac{1}{n}\sum_{k=1}^{n}\tilde{\sigma}^{2}_{k}$, and define
$U_{n}:=\tilde{\sigma}^{2}_{n}-\sigma^{2},$
such that $\mathbbm{E}(U_{n}|\mathcal{F}_{n-1})=\mathbbm{E}\big{(}\mathbbm{E}(U_{n}|\mathcal{F}_{\phi_{n-1}})\big{|}\mathcal{F}_{n-1}\big{)}=0$ and $\mathbbm{E}|U_{n}|\leq 2\sigma^{2}$, while
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}\big{(}U^{2}_{n}\big{|}\mathcal{F}_{n-1}\big{)}=\sum_{n=1}^{\infty}\frac{1}{n^{2}}\Big{(}\mathbbm{E}\big{(}\tilde{\sigma}^{4}_{n}\big{|}\mathcal{F}_{n-1}\big{)}-\sigma^{4}\Big{)}<\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}\big{(}\tilde{\sigma}^{4}_{n}\big{|}\mathcal{F}_{n-1}\big{)}.$
By using the expansion for the fourth central moment of an i.i.d. sum and using (34), we can find that
$\mathbbm{E}\big{(}\tilde{\sigma}^{4}_{n}\big{|}\mathcal{F}_{n-1}\big{)}$ $=\mathbbm{E}\Bigg{(}\frac{\mathbbm{E}\big{(}(Z_{n}-m\cdot\phi(Z_{n-1}))^{4}\big{|}\mathcal{F}_{\phi_{n-1}}\big{)}}{\phi^{2}(Z_{n-1})}\Bigg{|}\mathcal{F}_{n-1}\Bigg{)}$
$=\mathbbm{E}\Bigg{(}\frac{\mathbbm{E}(\xi-m)^{4}-3\sigma^{4}}{\phi(Z_{n-1})}+3\sigma^{4}\Bigg{|}\mathcal{F}_{n-1}\Bigg{)}$
$\leq\mathbbm{E}(\xi-m)^{4}.$
Since we have assumed that $\mathbbm{E}(\xi-m)^{4}$ is finite, it follows that
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}\big{(}U^{2}_{n}\big{|}\mathcal{F}_{n-1}\big{)}\leq\sum_{n=1}^{\infty}\frac{\mathbbm{E}(\xi-m)^{4}}{n^{2}}<\infty.$
Therefore, by Corollary 6.2, $\frac{1}{n}\sum_{k=1}^{n}U_{k}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\bar{\sigma}^{2}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma^{2}$.
(iv) Weak consistency of $\boldsymbol{\hat{\sigma}^{2}_{n}}$: We abandon the assumption that $m$ is known. We can decompose
$\hat{\sigma}^{2}_{n}=\bar{\sigma}^{2}_{n}+\underbrace{\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}(Z_{k}-m\cdot\phi(Z_{k-1}))}_{\text{I}}+\underbrace{\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\phi(Z_{k-1})}_{\text{II}},$
where we have shown in (iii) that $\bar{\sigma}^{2}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\sigma^{2}$. It remains to show that I and II both converge in probability to zero.
@@
    We can decompose
    $\hskip-17.07182pt\frac{2(m-\tilde{m}_{n})}{n}\sum_{k=1}^{n}(Z_{k}-m\cdot\phi(Z_{k-1}))=-2\underbrace{\sqrt{\frac{\sum_{k=1}^{n}Z_{k-1}}{\phi(Z_{n-1})}}}_{\text{(a)}}\cdot\underbrace{\frac{Z_{n}-m\cdot\phi(Z_{n-1})}{n^{1/4}\cdot\sqrt{\phi(Z_{n-1})}}}_{\text{(b)}}\cdot\underbrace{\frac{\sum_{k=1}^{n}(Z_{k}-m\cdot\phi(Z_{k-1}))}{n^{3/4}\sqrt{\sum_{k=1}^{n}Z_{k-1}}}}_{\text{(c)}},$
    and consider each of the components $(a)$, $(b)$ and $(c)$ in turn.
    1. (a)
      As a result of Lemma 6.4, for any $s$ such that $1<s<m\alpha$, we have that
      $\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sqrt{\frac{\sum_{k=1}^{n}Z_{k-1}}{Z_{n-1}}}\leq\sqrt{\frac{1}{1-s^{-1}}}\Bigg{)}=1.$
      Using Chebyshev’s inequality, for $r<\alpha$,
      $\mathbbm{P}(\phi(Z_{n-1})>r\cdot Z_{n-1}\,|\,Z_{n-1})$ $\geq 1-\mathbbm{P}\big{(}|\phi(Z_{n-1})-\alpha\cdot Z_{n-1}|\geq(\alpha-r)\cdot Z_{n-1}\,\big{|}\,Z_{n-1}\big{)}$
      $\geq 1-\frac{\beta}{(\alpha-r)^{2}\cdot Z_{n-1}}$
      By conditioning on the event $\{\phi(Z_{n-1})>r\cdot Z_{n-1}\}$, we then find that
      $\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sqrt{\frac{\sum_{k=1}^{n}Z_{k-1}}{\phi(Z_{n-1})}}\leq\sqrt{\frac{r}{1-s^{-1}}}\Bigg{)}$
      $=\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sqrt{\frac{\sum_{k=1}^{n}Z_{k-1}}{\phi(Z_{n-1})}}\leq\sqrt{\frac{r}{1-s^{-1}}}\,\Bigg{|}\,\phi(Z_{n-1})>r\cdot Z_{n-1}\Bigg{)}\cdot\mathbbm{P}(\phi(Z_{n-1})>r\cdot Z_{n-1})$
      $=1$
      on the event $\{Z_{n}\to\infty\}$. Hence we see that $\sqrt{\frac{\sum_{k=1}^{n}Z_{k-1}}{\phi(Z_{n-1})}}$ is asymptotically bounded by a constant in probability.
    2. (b)
      Given that
      $\mathbbm{E}\bigg{(}\frac{Z_{n}-m\cdot\phi(Z_{n-1})}{n^{1/4}\cdot\sqrt{\phi(Z_{n-1})}}\bigg{|}\mathcal{F}_{\phi_{n-1}}\bigg{)}=0\quad\text{and}\quad\text{Var}\bigg{(}\frac{Z_{n}-m\cdot\phi(Z_{n-1})}{n^{1/4}\cdot\sqrt{\phi(Z_{n-1})}}\bigg{|}\mathcal{F}_{\phi_{n-1}}\bigg{)}=\frac{\sigma^{2}}{\sqrt{n}},$
      we see that
      $\mathbbm{E}\bigg{(}\frac{Z_{n}-m\cdot\phi(Z_{n-1})}{n^{1/4}\cdot\sqrt{\phi(Z_{n-1})}}\bigg{)}=0\quad\text{and}\quad\text{Var}\bigg{(}\frac{Z_{n}-m\cdot\phi(Z_{n-1})}{n^{1/4}\cdot\sqrt{\phi(Z_{n-1})}}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
      so, as a consequence of Chebyshev’s inequality, it follows that
      $\frac{Z_{n}-m\cdot\phi(Z_{n-1})}{n^{1/4}\cdot\sqrt{\phi(Z_{n-1})}}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0.$
    3. (c)
      Defining
      $U_{n}:=Z_{n}-m\cdot\phi(Z_{n-1})\quad\text{and}\quad J_{n}:=n^{3/4}\sqrt{\sum_{k=1}^{n}Z_{k-1}},$
      we can see that $\{U_{n}\}_{n\in\mathbbm{N}_{0}}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{n}\}_{n\in\mathbbm{N}_{0}}$, since $\mathbbm{E}(U_{n}|\mathcal{F}_{n-1})=0$ for all $n\in\mathbbm{N}_{1}$ and
      $\mathbbm{E}|U_{n}|$ $\leq\sqrt{\mathbbm{E}(U_{n}^{2})}$ (Hölder’s inequality)
      $=\sqrt{\mathbbm{E}(\text{Var}(Z_{n}|\mathcal{F}_{\phi{n-1}}))}$
      $=\sqrt{\sigma^{2}\cdot\mathbbm{E}(\phi(Z_{n-1}))}$
      $=\sqrt{\sigma^{2}\alpha\cdot\mathbbm{E}(Z_{n-1})}$
      $<\infty,$
      and $\{J_{n}\}_{n\in\mathbbm{N}_{0}}$ is positive, non-decreasing, with $\lim_{n\to\infty}J_{n}=\infty$. Additionally,
      $\sum_{n=1}^{\infty}J_{n}^{-2}\mathbbm{E}(U_{n}^{2}|\mathcal{F}_{n-1})$ $=\sum_{n=1}^{\infty}\frac{\mathbbm{E}(\text{Var}(Z_{n}|\mathcal{F}_{\phi_{n-1}})|\mathcal{F}_{n-1})}{n^{3/2}\cdot\sum_{k=1}^{n}Z_{k-1}}$
      $=\sum_{n=1}^{\infty}\frac{\sigma^{2}\alpha\cdot Z_{n-1}}{n^{3/2}\cdot\sum_{k=1}^{n}Z_{k-1}}$
      $<\sum_{n=1}^{\infty}\frac{\sigma^{2}\alpha}{n^{3/2}}$
      $<\infty\;\;\text{a.s.},$
      so Theorem 6.1 applies, such that $J_{n}^{-1}\sum_{k=1}^{n}U_{k}=\frac{\sum_{k=1}^{n}(Z_{k}-m\cdot\phi(Z_{k-1}))}{n^{3/4}\sqrt{\sum_{k=1}^{n}Z_{k-1}}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ as $n\to\infty$.